OpenStudy (anonymous):

\(\sqrt{48}\) , \(sqrt{96}\) Simplify ty :)

OpenStudy (anonymous):

\(\sqrt{96}\) ***

OpenStudy (asnaseer):

you need to use exactly the same method here. basically you need to transform the number under the squareroot sign into a product involving perfect squares. so try to get it into the form:\[\sqrt{a^2b}=\sqrt{a^2}*\sqrt{b}=a*\sqrt{b}=a\sqrt{b}\]

OpenStudy (anonymous):

simplify each or are you multiplying them?

OpenStudy (anonymous):

Simplify each

OpenStudy (asnaseer):

so how can you write 48 as the product of a perfect square and something else?

OpenStudy (asnaseer):

\[\begin{align} 4=2^2\\ 9=3^2\\ 16=4^2\\ 25=5^2\\ 36=6^2\\ \end{align}\]

OpenStudy (anonymous):

asnaseer is trying to teach you how to do the work if i write the answer il ruin it for you try to figure it out with his examples if you really dont il post the answer later

OpenStudy (asnaseer):

thank for that suzy4321 - much appreciated :)

OpenStudy (anonymous):

sorry took dog out, let me read this :)

OpenStudy (anonymous):

\[4\sqrt{3}\]?

OpenStudy (asnaseer):

perfect! - you're a pro now! :)

OpenStudy (asnaseer):

you should be able to do the same with \(\sqrt{96}\) now

OpenStudy (asnaseer):

notice that 96 = 48 * 2

OpenStudy (anonymous):

\[4\sqrt{6}

OpenStudy (anonymous):

\[4\sqrt{6}

OpenStudy (asnaseer):

apart from the latex - your answer is correct again - well done!

OpenStudy (anonymous):

\[4\sqrt{6}\]

OpenStudy (asnaseer):

yup - you got it

OpenStudy (anonymous):

kk ill finish these then get back to you, cus it gets harder soon, if your still on :)

OpenStudy (asnaseer):

its past midnight here and I'm off to catch some Zzzz's now - but good luck and I'm sure there'll be plenty of other people willing to help you here.

OpenStudy (anonymous):

ahh okay than